OpenStudy (anonymous):

Which of the following is an actual zero of f(x) = x3 + x2 - 4x - 4? (4 points) Question 4 options: 1) 4 2) 1 3) -1 4) -4

OpenStudy (solomonzelman):

For this question you don't need anything beyond algebra that is usually applied in college algebra. Just factoring. \(\normalsize\color{black}{ f(x) = x^3 + x^2 - 4x - 4}\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ f(x) = x^2(x + 1) - 4(x +1)}\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ f(x) =( x^2-4)(x + 1) }\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ f(x) =(x+2)( x-2)(x + 1) }\)\(\LARGE\color{white}{ \rm │ }\) do you need more help? let me know if you do.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

(x-2)(x+2) cancel right

OpenStudy (anonymous):

i got x = -1

OpenStudy (anonymous):

how

OpenStudy (anonymous):

idk. i plugged it into solomen zelmens thingy lol

OpenStudy (anonymous):

@SolomonZelman im confused lol

OpenStudy (anonymous):

did you cancel out the 2's

OpenStudy (anonymous):

its not 1 bc 1 ^3 is 1 and 1^ is 1 and 4x1 is 4 so... 1+1-4-4= -6 not 0

OpenStudy (anonymous):

idk i give up

OpenStudy (solomonzelman):

I am sorry for late reply, but when you have an equation such as \(\normalsize\color{black}{ f(g)=(x+v)(x-a)(x+s)}\) the roots are going to be \(\normalsize\color{black}{ -v,~~~~a~~~~~and~~~-s}\)

OpenStudy (solomonzelman):

So, for \(\normalsize\color{black}{ f(g)=(x+2)(x-2)(x+1)}\) \(\LARGE\color{white}{ \rm │ }\) the roots are going to be equal to \(\normalsize\color{black}{ 2,-2,-1}\) . \(\LARGE\color{white}{ \rm │ }\) Let me know ifthere is anything that's troubling you.

OpenStudy (anonymous):

ohhh okay I see